space vehicles use solid lithium hydroxide to remove exhaled carbon dioxide according to the equation 2lioh +co2 -> li2co3 +h2o determine the mass of carbon dioxide removed if the space vehicle carries 42.0 mol lioh


1
Expert's answer
2021-12-06T13:50:02-0500

Mole ratio of LiOH: CO2=2:1

= 42/2= 21moles

Molar mass of CO2= 44.01

= 21×44.01=924.21g